OpenStudy (anonymous):

find dy/dx of 3xy^2 = 5y + 2

OpenStudy (anonymous):

dy/dx=-(3y^2)/(6xy-5)

OpenStudy (anonymous):

\[6xy(dy) + 3y^{2}(dx) = 5dy\]\[\frac{dy}{dx} = \frac{-3y^{2}}{6xy - 5}\]